OpenStudy (anonymous):

A recipe for potting soil that makes enough to fill 3 bags calls for 5 1/2 lb of compost. David wants to modify the recipe so that it fills 5 bags. How many more pounds of compost will David need for the modified recipe than the original recipe calls for?

OpenStudy (anonymous):

is there any more?

OpenStudy (anonymous):

Any more questions that I need help with?

OpenStudy (anonymous):

no any more to the problem?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Well There are the answers : A- \[3 \frac{ 2 }{ 3 }\] B- \[6 \frac{ 3 }{ 5 }\] C- \[9 \frac{ 1 }{ 6 }\] D- 27 \[27 \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

that helps alot lol

OpenStudy (anonymous):

@gg54